# QAEA-DR: A Unified Text Augmentation Framework for Dense Retrieval

### v1

@@ II-A Dense Retrieval
Dense retrieval has become an important research area following the development of pre-trained Transformer language models (PLMs) [19, 20, 2, 21, 22]. To enhance text retrieval performance, dense retrieval leverages PLM-based text embeddings to encode queries and documents into a shared semantic vector space, focusing on matching semantic contents beyond mere keywords. This text embedding application in retrieval is fundamental to Retrieval-Augmented Generation (RAG) [23, 24], which reduces the hallucinations in LLMs. Recent advancements in dense retrieval include architectural innovations, optimized training methodologies, and efficient indexing techniques, all of which contribute to improved retrieval accuracy and efficiency [4, 5, 6, 7, 8, 25]. Since the introduction of Sentence-BERT [26] and Dense Passage Retrieval (DPR) [2], numerous sentence embedding models have been proposed to enhance dense passage retrieval. Advanced sentence embedding models, which have been highlighted in the retrieval task of massive text embedding benchmark (MTEB) [27], include Contriever [28], M3E [^2] https://huggingface.co/moka-ai/m3e-base, BGE [9], etc. Our text augmentation method serves as a preprocessing module for dense retrieval and is compatible with various embedding models mentioned above.
@@ II-B Input Enhancement in Retrieval
In addition to the optimization methods for the retriever, input enhancement strategy represents a distinct optimization approach for retrieval tasks [10]. In particular, input data to a retriever includes user query and datastore. Therefore, input enhancement for retrieval can be categorized into two types: query transformation and data augmentation. Query transformation modifies the input query during retrieval, for example, Hypothetical Document Embeddings (HyDE) [12] that generate pseudo documents from queries, and KNN-based Query Expansion (KNN-QE) [11] that enhances queries using local conceptual word embeddings. Data augmentation improves the data to be retrieved before the retrieval process, including synthesizing data, clarifying ambiguities, updating outdated data, etc. Compared to query transformation, data augmentation models have the advantage of not consuming user waiting time in retrieval, which is particularly important in practical applications. Mainstream studies focus on data augmentation for cross-modal retrieval, such as Make-An-Audio [29], AMSH [30], and ReACC [31]. In terms of text-to-text retrieval, methods like InPars [13] generates new query-text pairs as training data. As current data augmentation methods do not consider enhancing the original text in text retrieval, we propose a text augmentation framework in this paper.
@@ II-C Information Extraction
  • EE has evolved from rule-based approaches [35] to machine learning methods like Dynamic Multi-Pooling Convolutional Neural Networks (DMCNN) [36] and Joint Event Extraction via Recurrent Neural Networks (JRNN) [37], and more recently to ChatGPT for Event Extraction (ChatEE) [38], and the Generative Template-based Event Extraction (GREE) [39], reflecting significant progress in the field.
@@
  • QAG have progressed from rule-based models [44, 45, 46] to generative-based PLMs like Information-Maximizing Hierarchical Conditional VAEs (Info-HCVAE) [47] and Language Models for Question and Answer Generation (LMQG) [48].
@@ IV-A Datasets and Baselines
Datasets. We utilize four passage retrieval datasets to evaluate our QAEA-DR. Due to the high computational cost of multiple LLM tasks, we used a subset of the complete open datasets for our experiments.
@@ IV-B Implementation Details
LLM Implementation. We employ two LLMs as text augmentation generators and another LLM as a evaluator for evaluating the quality of the generated texts. In our experiments, we utilize API calls to GPT-3.5-turbo and locally deployed ChatGLM3-6B [53] as the two generators for QAG and EE. We set the temperature parameter to 0 to ensure the generated text is precise and faithful to the original input. In our experiments, since ChatGLM3-6B is not friendly with JSON output, we manually create 20 JSON samples for fine-tuning to ensure the stable JSON output. Additionally, we select DeepSeek-v2333https://platform.deepseek.com as the evaluator for scoring the quality of generated texts, using API calls for text generation. Given that DeepSeek-v2 significantly outperforms GPT-3.5-turbo and ChatGLM3-6B on LLM evaluation benchmarks, it serves as an effective independent evaluator to enhance the reliability of the output assessments. All experiments run on the NVIDIA RTX 3090 GPU (24 GB).
@@ IV-E Ablation Study on QAEA
Figure 4: Ablation Performance of QAEA: NDCG@1 (×$\times 100$)
@@
V Analysis
In this section, we conduct the analysis of QAEA-DR framework and evaluate its performance under different conditions and configurations.
@@ V-B Impact of Regeneration Threshold
TABLE V: Average Number of QA Pairs and Events Generated From Original Text Through LLMs
@@ V-C Impact of Generated Text Quantity
We investigate the impact of the number of generated texts per original text on QAEA-DR performance. Table V summarizes the average number of QA pairs and events generated from original texts using GPT or GLM across four datasets. In all datasets, GPT generates more structured texts than GLM. Referring to Table I, where GPT generally outperforms GLM in QAEA (TRI) retrieval performance, we infer a positive correlation between the number of generated texts and retrieval performance. Table V also indicates that the average number of QA pairs generated is 2.15 times the average number of events. This aligns with intuition since each QA pair only conveys an individual information point, while event summarizes richer information.
TABLE VI: Comparison of LMQG and Our LLM-based GPT-QAG on Retrieval Performance: NDCG@1 (×$\times 100$)
@@
V-D QAG Comparisons
Table VI presents the QAG retrieval performance of our LLM-based text augmentation method compared to the popular QAG model LMQG. We use the QA component of the QAEA framework as described in Section IV-E and employ GPT-3.5-turbo as the text generation LLM, denoted as GPT-QAG. Both LMQG and GPT-QAG generate vector databases of QA pair texts $\text{VDB}_{\text{QA}}$ for dense retrieval. Since LMQG only supports English, we display comparison results on two English datasets. We observe that under both TRI and TMO QA text organization strategies, our GPT-QAG significantly outperforms LMQG, demonstrating the effectiveness of LLM-based QAG using a three-step prompting process. Notably, evaluating the text retrieval capability of generated QA pairs offers a new insight for assessing QAG performance.
TABLE VII: Time Complexity Comparisons for QAEA (TMO) and QAEA (TRI) on sCNPR
@@ V-D QAG Comparisons
V-E Time Complexity
The computational time of dense retrieval is independent of text length since all texts are embedded into vectors of the same dimensionality. Hence, the retrieval time is directly proportional to the number of text vectors in the vector database. For $n$ texts and the associated vector database $\text{VDB}_{\text{ori}}=\{v_{1},v_{2},\ldots,v_{n}\}$ under baseline, the time taken for each query to traverse the texts is $n\times T_{\text{sim}}$, where $T_{\text{sim}}$ is the computational cost to calculate $\text{sim}(v_{q},v_{i})$, $v_{i}\in\text{VDB}_{\text{ori}}$. Table VII presents the time complexity for QAEA (TMO) and QAEA (TRI) on the sCNPR dataset. Similar conclusions regarding time complexity are observed across other datasets.

### v2

@@ II-A Dense Retrieval
Dense retrieval has become an important research area following the development of pre-trained Transformer language models (PLMs) [19, 20, 2, 21, 22]. To enhance text retrieval performance, dense retrieval leverages PLM-based text embeddings to encode queries and documents into a shared semantic vector space, focusing on matching semantic contents beyond mere keywords. This text embedding application in retrieval is fundamental to Retrieval-Augmented Generation (RAG) [23], which reduces the hallucinations in LLMs. Recent advancements in dense retrieval include architectural innovations, optimized training methodologies, and efficient indexing techniques, all of which contribute to improved retrieval accuracy and efficiency [4, 5, 6, 7, 8, 24]. Since the introduction of Sentence-BERT [25] and Dense Passage Retrieval (DPR) [2], numerous sentence embedding models have been proposed to enhance dense passage retrieval. Advanced sentence embedding models, which have been highlighted in the retrieval task of massive text embedding benchmark (MTEB) [26], include Contriever [27], M3E [^2] https://huggingface.co/moka-ai/m3e-base, BGE [9], etc. Our text augmentation method serves as a preprocessing module for dense retrieval and is compatible with various embedding models mentioned above.
@@ II-B Input Enhancement in Retrieval
In addition to the optimization methods for the retriever, input enhancement strategy represents a distinct optimization approach for retrieval tasks [10]. In particular, input data to a retriever includes user query and datastore. Therefore, input enhancement for retrieval can be categorized into two types: query transformation and data augmentation. Query transformation modifies the input query during retrieval, for example, Hypothetical Document Embeddings (HyDE) [12] that generate pseudo documents from queries, and KNN-based Query Expansion (KNN-QE) [11] that enhances queries using local conceptual word embeddings. Data augmentation improves the data to be retrieved before the retrieval process, including synthesizing data, clarifying ambiguities, updating outdated data, etc. Compared to query transformation, data augmentation models have the advantage of not consuming user waiting time in retrieval, which is particularly important in practical applications. Mainstream studies focus on data augmentation for cross-modal retrieval, such as Make-An-Audio [28] and ReACC [29]. In terms of text-to-text retrieval, methods like InPars [13] generates new query-text pairs as training data. As current data augmentation methods do not consider enhancing the original text in text retrieval, we propose a text augmentation framework in this paper.
@@ II-C Information Extraction
  • EE has evolved from rule-based approaches [33] to deep learning methods like Dynamic Multi-Pooling Convolutional Neural Networks (DMCNN) [34] and Joint Event Extraction via Recurrent Neural Networks (JRNN) [35], and more recently to ChatGPT for Event Extraction (ChatEE) [36], and the Generative Template-based Event Extraction (GREE) [37], reflecting significant progress in the field.
@@
  • QAG have progressed from rule-based models [42] to generative-based PLMs like Information-Maximizing Hierarchical Conditional VAEs (Info-HCVAE) [43] and Language Models for Question and Answer Generation (LMQG) [44].
@@ IV-A Datasets and Baselines
Datasets. We utilize four passage retrieval datasets to evaluate our QAEA-DR. Due to the high computational cost of multiple LLM tasks, we used a subset of the complete open datasets for our experiments. To ensure the unbiased selection of dataset subsets, we applied the random.sample() method in Python to randomly select texts from the corpus.
@@ IV-B Implementation Details
LLM Implementation. We primarily employ two LLMs as text augmentation generators and another LLM as a evaluator for evaluating the quality of the generated texts. In our experiments, we utilize API calls to GPT-3.5-turbo and locally deployed ChatGLM3-6B [49] as the two generators for QAG and EE. Additionally, in Section IV-G, we deployed Mistral-7B [50] and Llama-3-8B [51] to compare the performance of different open-source LLMs. We set the temperature parameter to 0 to ensure the generated text is precise and faithful to the original input. In our experiments, since ChatGLM3-6B is not friendly with JSON output, we manually create 20 JSON samples for fine-tuning to ensure the stable JSON output. Additionally, we select DeepSeek-v2333https://platform.deepseek.com as the evaluator for scoring the quality of generated texts, using API calls for text generation. Given that DeepSeek-v2 significantly outperforms GPT-3.5-turbo and ChatGLM3-6B on LLM evaluation benchmarks, it serves as an effective independent evaluator to enhance the reliability of the output assessments. All experiments run on the NVIDIA RTX 3090 GPU (24 GB).
@@ IV-E Ablation Study on QAEA
Figure 4: Ablation Performance of QAEA: NDCG@1 (×$\times 100$)
@@
TABLE V: The NDCG (×$\times 100$) performance comparison of different LLMs with QAEA (TMO)
Models HotpotQA MS MARCO
dpr bge-en dpr bge-en
NDCG@1 NDCG@10 NDCG@1 NDCG@10 NDCG@1 NDCG@10 NDCG@1 NDCG@10
Baseline 79.38 74.24 95.79 92.70 70.23 80.64 95.93 98.02
GPT-3.5-turbo 81.97 76.01 96.06 92.91 74.33 83.80 96.21 98.19
ChatGLM3-6B 81.66 75.96 95.99 92.83 74.96 83.72 96.17 98.15
[…]
$\dagger$ denotes zero-shot prompting without optimization through evaluation and regeneration mechanisms.
IV-F Case Study
To better show how generated QA pairs and events address key information loss in long and noisy texts, we present the outputs of QAEA-DR via GPT-3.5-turbo on two cases from the MS MARCO dataset, as shown in Table VI. Table VI displays the generated texts with the highest score for each case.
[…]
IV-G More LLMs Generation Results
For further experiments on QAEA-DR with more open-source LLMs as text augmentation generators, we adopt Mistral-7B [50] and Llama-3-8B [51]. Both Mistral and Llama use standard zero-shot prompts without evaluation and regeneration steps, while GPT-3.5-turbo and ChatGLM3-6B apply scoring-based evaluation and regeneration to improve output quality.
The results in Table V show that GPT achieves the highest NDCG@1, especially with bge-en. GLM is slightly lower but still competitive. Mistral and Llama perform slightly weaker than GLM with regeneration. However, Mistral and Llama are likely to improve if they also support regeneration. Among open-source LLMs, ChatGLM3-6B is a strong choice for QAEA-DR. It supports both Chinese and English and has a smaller parameter size. All three open-source LLMs outperform the baselines and show performance close to GPT-3.5-turbo on English datasets. These findings confirm that small-scale open-source LLMs can provide effective text augmentation for QAEA-DR.
TABLE VI: Case study on QAEA-DR
Case1 Myers-briggs personality test what do the letters stand for?
Gold Original Text
{“text”: “INTP (introversion, intuition, thinking, perceiving)
[…]
MBTI’s 16 personality types. NTP (introversion, intuition, thinking,
perceiving) is an abbreviation used in the publications of the
[…]
of the MBTI’s 16 personality types.”}
Score: 37.73           Rank: 324
QA Pair
{“text”: “What does INTP stand for in the context of the
Myers-Briggs Type Indicator (MBTI)? INTP stands for introversion,
intuition, thinking, and perceiving, and it refers to one of the 16
personality types in the MBTI.”}
Score: 72.59           Rank: 1
Event
{“text”: “Definition INTP abbreviation used in the publications
of the Myers Briggs Type Indicator (MBTI) to refer to one of the
MBTI’s 16 personality types.”}
Score: 61.08           Rank: 4
Case2 How much will i need to close on a house with fha?
Gold Original Text
{“text”: “And then there are closing costs. Those two costs will
equal the total cash needed to close. Closing costs are roughly 2-2.25%
of the purchase price. So if you’re looking at a home for \$500,000, plan on
closing costs to be around \$10,000.How much down payment … ” }
(160 words)
Score: 36.32           Rank: 135
QA Pair
{“text”: “What are closing costs and how are they calculated? Closing
costs are one of the two costs that equal the total cash needed to
close a property purchase. They are roughly 2-2.25% of the purchase price.”}
Score: 73.33           Rank: 2
Event
{“text”: “Closing Costs. Closing costs. purchase price. Closing
costs are roughly 2-2.25% of the purchase price. For a home priced at
\$500,000, closing costs would be around \$10,000.”}
Score: 76.04           Rank: 1
V Analysis
In this section, we conduct the analysis of QAEA-DR and evaluate its performance under different conditions and configurations.
@@ V-B Impact of Regeneration Threshold
TABLE VII: Average Number of QA Pairs and Events Generated From Original Text Through LLMs
@@ V-C Impact of Generated Text Quantity
We investigate the impact of the number of generated texts per original text on QAEA-DR performance. Table VII summarizes the average number of QA pairs and events generated from original texts using GPT or GLM across four datasets. In all datasets, GPT generates more structured texts than GLM. Referring to Table I, where GPT generally outperforms GLM in QAEA (TRI) retrieval performance, we infer a positive correlation between the number of generated texts and retrieval performance. Table VII also indicates that the average number of QA pairs generated is 2.15 times the average number of events. This finding is intuitive since each QA pair only conveys an individual information point, while event summarizes richer information.
TABLE VIII: Comparison of QAEA-DR generated text diversity scores
Metrics MS MARCO HotpotQA
QA Event QA Event
GPT GLM GPT GLM GPT GLM GPT GLM
Compression Ratio 1.0896 1.1016 1.2242 1.6411 0.9991 1.0112 1.1057 1.5182
Self-BLEU 0.0934 0.0158 0.1457 0.4000 0.0414 0.0088 0.0867 0.3776
Self-BERTScore 0.6848 0.6806 0.6662 0.8606 0.6355 0.6364 0.6006 0.8482
Self-Repetition 0.1524 0.2024 0.0875 0.4786 0.0986 0.1623 0.0462 0.4595
NDCG@1 71.67 71.78 65.29 64.80 79.07 79.27 73.01 72.24
Lower diversity scores suggest greater text diversity.
V-D Impact of Generated Text Diversity
We perform a generated text diversity analysis to measure how the diversity of QA pairs and events affects retrieval performance, as shown in Table VIII. The analysis includes several metrics: Compression Ratio, Self-BLEU, Self-BERTScore, and Self-Repetition, where lower scores suggest greater text diversity [52].
The results show a clear correlation between higher diversity (lower scores) and improved retrieval performance. This effect is particularly evident in generated events across two datasets. Specifically, QA pairs exhibit greater diversity and better retrieval performance than events. For example, QA pairs have an average Self-BLEU score of 0.0399 with an NDCG@1 score of 75.35, while events have a higher Self-BLEU score of 0.2525 and an NDCG@1 score of 68.84. Additionally, although GPT generally shows more diversity than GLM across the two English datasets, their NDCG@1 scores remain comparable. This suggests that minor variations in text diversity have limited influence on retrieval performance, highlighting the robustness of QAEA-DR. We also observe that Self-BLEU strongly correlates with retrieval performance. Self-BLEU serves as a valuable metric for evaluating generated text quality, offering a lexical perspective alongside embeddings.
TABLE IX: Comparison of LMQG and Our LLM-based GPT-QAG on Retrieval Performance: NDCG@1 (×$\times 100$)
@@ V-D Impact of Generated Text Diversity
V-E QAG Comparisons
Table IX presents the QAG retrieval performance of our LLM-based text augmentation method compared to the popular QAG model LMQG. We use the QA component of the QAEA framework as described in Section IV-E and employ GPT-3.5-turbo as the text generation LLM, denoted as GPT-QAG. Both LMQG and GPT-QAG generate vector databases of QA pair texts $\text{VDB}_{\text{QA}}$ for dense retrieval. Since LMQG only supports English, we display comparison results on two English datasets. We observe that under both TRI and TMO QA text organization strategies, our GPT-QAG significantly outperforms LMQG, demonstrating the effectiveness of LLM-based QAG using a three-step prompting process. Notably, evaluating the text retrieval capability of generated QA pairs offers a new insight for assessing QAG performance.
TABLE X: The NDCG (×$\times 100$) and generated text noise (%) performance on MS MARCO dataset under different input noise levels (%)
Method Input Noise (%) NDCG@1 Generated Text Noise (%)
QAEA-DR 0 96.21
20 96.12 0.33
50 96.15 0.80
100 96.15 1.94
QA 0 93.98
20 94.34 0.05
50 94.00 0.13
100 92.18 0.20
Event 0 91.56
20 92.12 0.79
50 91.78 1.86
100 91.86 4.38
V-F Text Noise Sensitive Analysis
We conduct a sensitivity analysis to evaluate the robustness of QAEA-DR against noise. We note that the noise in the texts refers to content that is unrelated to the query. In our experiments, we introduce semantically meaningless texts into the original texts to simulate random noise, defined as sequences of randomly generated letters, digits, and spaces (e.g., DtAfgJndLj FO7djxaOM). The original texts are augmented with 20%, 50%, and 100% input noise based on their text length and processed through QAEA-DR to generate QA and event texts. We compare the NDCG@1 scores for retrieval with and without noise and calculate the proportion of input noise retained in the generated texts.
Table X summarizes the results using GPT-3.5-turbo generator and the bge-en embedding model under the QAEA (TMO) strategy. The table shows that retrieval performance is not greatly affected by adding significant input noise. QA pairs and events retain some noise from the input, but the proportion of retained noise is much lower. For example, with 100% input noise, the generated QA pairs retain 0.20% of the noise, while the generated events retain 4.38%. The higher noise retention in events is likely due to their structured complexity and longer descriptions. QA pairs focus on individual information points, which makes them less likely to retain noise.
TABLE XI: Time Complexity Comparisons for QAEA (TMO) and QAEA (TRI) on sCNPR
@@ V-F Text Noise Sensitive Analysis
TABLE XII: Computational cost comparison of different LLMs on MS MARCO dataset
LLMs Task Size NDCG@1 Time Memory
GPT-3.5-turbo QAG Unknown 93.98 3.29s Unknown
EE 91.56 2.69s
ChatGLM3 QAG 6B 93.28 11.66s 11.9GB
EE 90.86 6.91s
Mistral QAG 7B 93.23 15.26s 13.8GB
EE 92.01 6.58s
Llama-3 QAG 8B 92.89 10.39s 15.3GB
EE 87.67 6.05s
V-G Retrieval Time Complexity
The computational time of dense retrieval is independent of text length since all texts are embedded into vectors of the same dimensionality. Hence, the retrieval time is directly proportional to the number of text vectors in the vector database. For $n$ texts and the associated vector database $\text{VDB}_{\text{ori}}=\{v_{1},v_{2},\ldots,v_{n}\}$ under baseline, the time taken for each query to traverse the texts is $n\times T_{\text{sim}}$, where $T_{\text{sim}}$ is the computational cost to calculate $\text{sim}(v_{q},v_{i})$, $v_{i}\in\text{VDB}_{\text{ori}}$. Table XI presents the time complexity for QAEA (TMO) and QAEA (TRI) on the sCNPR dataset. Similar conclusions regarding time complexity are observed across other datasets.
@@ V-G Retrieval Time Complexity
V-H LLM Computational Costs
Table XII compares the computational costs of generating QAEA-DR texts for GPT-3.5-turbo, ChatGLM3-6B, Mistral-7B, and Llama-3-8B. The NDCG@1 results in Table XII are based on the bge-en embedding model under the QAEA (TMO) strategy. All experiments run on the NVIDIA RTX 3090 GPU (24 GB) without optimizations for inference efficiency or memory usage.
We observe that QAG tasks require more inference time than EE tasks, which suggests that QAG outputs are generally longer. This difference in output characteristics may contribute to QA pairs achieving higher NDCG@1 scores compared to events. GPT-3.5-turbo outperforms open-source small models in both accuracy and inference time, demonstrating its superior efficiency. For open-source LLMs, memory usage scales with model size, yet smaller-scale models still achieve reasonable inference time and solid retrieval performance. These findings show that QAEA-DR can run on moderate hardware in resource-limited scenarios, while larger models can improve generation quality. This flexibility allows users to balance computational cost and retrieval performance.